# Beyond Silence: Bias Analysis through Loss and Asymmetric Approach in Audio Anti-Spoofing

### v1

@@ 1 Introduction
Recent progress in voice conversion (VC) and text-to-speech (TTS) technologies have intensified concerns regarding their potential for malicious use, emphasizing the critical role of audio anti-spoofing systems. Audio spoofing detection systems, as binary classifiers, distinguish genuine human speech (bonafide) from artificially generated (spoofed) speech. The main direction in this field is toward advancing the systems’ ability to generalize across unseen spoofing attacks by focusing on learning diverse spoofing artifacts. To this end, a pivotal shift from traditional hand-crafted features [1, 2] to data-driven approaches [3, 4, 5] and data augmentation [6, 7, 8] have been pursued. Research on diversifying training data [9, 10, 11], incorporating domain adaptation strategies [12, 13, 14], and leveraging large-scale pre-trained models [15, 16, 17] also follow this trend.
@@
While in-depth analyses in audio anti-spoofing have significantly advanced the field and provided invaluable insights, several areas remain for further explorations: (i) the majority of studies have focused on spoof class, specifically attack-wise interpretations; (ii) class-wise analyses exist, yet the emphasis has primarily been on silence, except for [30] exploring other data bias factors; (iii) existing analyses predominantly rely on evaluating outcomes, such as performance metrics and score distributions, without a thorough examination of the internal workings of models during training. Our research diverges from existing studies by delving into the training process through loss analysis and adopting a novel asymmetric approach for class/phase-wise interpretation.
Our analysis starts with examining the loss curves for bonafide and spoof classes separately, both with and without the application of RawBoost [7] data augmentation, as illustrated in Figure 1. Note that illustrated loss values are raw values. For the model update, a loss weight of 0.9 (for bonafide) and 0.1 (for spoof) was employed, considering the imbalanced number of samples in each class. Interestingly, the results indicate that the loss associated with the bonafide class is significantly lower than that for the spoof class. Even if one considers the loss weights in the training phase, a substantial gap still remains between the two classes. These results could signify that the bonafide class is inherently easier to train than the spoof class. Alternatively, this might suggest that modeling the bona fide class is not trivial; however, there exists a shortcut that significantly reduces the training loss. We deploy various loss functions to reveal the meaning behind the low magnitude of the bonafide loss, between the two scenarios mentioned. In particular, the objective functions that can consider the difficulty of samples enable us to understand how the model deals with each class based on our findings.
@@ 3.1 Datasets
ASVspoof2019 [27] is a dataset dedicated to developing audio anti-spoofing and spoofing-robust automatic speaker verification systems. It includes 6 and 13 types of spoofing attacks in train/dev and test. The number of utterances for train/dev and test is $50,224$ and $71,237$, respectively.
@@
ASVspoof2021 [39] includes the latest spoofing attacks, which consists of different test sets, each for LA and DF scenarios. The attacks in the LA subset are processed with real telephony systems introducing artifacts related to encoding and transmission. The attacks in the DF subset includes various audio compressions.
@@ 3.2 Implementation details
Table 3: Performance comparison when the model is tested on asymmetric ways. All models are trained on the original training dataset but evaluated in different configurations. Here, the ratio indicates a relative change of Te_B compared to that of Te_S.
@@
Table 5: Performance comparison in EER when the model trained on asymmetric ways. All evaluations are conducted using original test set without interventions. The bold type indicates the case that spoof side intervention lower EER than the bonafide one.
@@ 4.2 Asymmetric test results
Table 3 presents the results of applying asymmetric interventions to a specific class during the test phase. These results are analyzed from two angles: (i) evaluating the intervention’s impact by comparing the original results (O) with those after interventions on the spoof (Te_S) or bonafide (Te_B) classes (columns 2 and 3), and (ii) examining how interventions differently affect the bonafide class compared to the spoof class, as indicated in column 4. From the first angle, enhancements in performance post-intervention hint at possible model biases due to shortcuts. Conversely, declines indicate the intervention’s irrelevance to the target class or might reflect domain differences or unknown factors. The “Ratio” in column 4, calculated as the relative performance impact $(|\textbf{O}-\textbf{Te}\_\textbf{B}|/\textbf{O})/(|\textbf{O}-\textbf{Te}\_\textbf{S}|/\textbf{O})$, offers deeper insight into which class the intervention impacts more. Here, $\text{Ratio}>1$ refers to a greater influence on the bonafide class, while $\text{Ratio}<1$ implies spoof class is more affected.
Our findings highlight two main observations that align with our analysis of loss functions. First, interventions are prone to improve Te_S (e.g., AASIST with MP3 or AASIST-L with Noise on Te_S), while consistently lowering Te_B performance. This implies that factors leading to quality degradation in utterances tend to be linked with the spoof class, helping the model to classify such inputs as spoofed. The drop in Te_B performance may result from interventions causing bonafide utterances to resemble spoofed ones more closely. Second, the “Ratio” always surpasses 1, as a universal trend regardless of which intervention applied. This emphasizes the bonafide class’s modeling fragility, suggesting that future research should more focus on the robust modeling of the bonafide class.
@@ 4.3 Further test interventions with silence trimming
As introduced in Section 1, the silence could significantly influence the class-wise imbalance results. This prompts an important question: “Do our observations remain valid when silence is excluded from the training or testing datasets or both?” To address this, we conducted experiments with silence removed – referred to as silence trimming – to assess its effect on our findings and show the results in  Table 4. The results show a remarkable consistency across most of our results. When silence was eliminated from both phases, effectively eliminating its influence, the ratio consistently exceeded 1, reaching a peak of 46.09. In additional experiments where silence was removed from only one phase, there were a few instances where the ratio fell below 1; however, the majority still exhibited ratios significantly greater than 1. These consistent outcomes robustly support the conclusion that the presence of silence does not skew our findings.
@@ 4.4 Asymmetric training results
Lastly, we focus on asymmetric interventions during the training phase to explore two key questions: (i) “What happens when interventions are applied solely during training?” and (ii) “Does the class-specific effect of interventions remain consistent if the test set remains untouched?” [^2] The second question is related to the analysis in Section 4.2, where we want to clarify whether performance degradation stems from suddenly appearing unseen domains in the test phase. According to the findings presented in Table 5, an improvement was observed in only one out of nine instances for the Tr_S condition. In contrast, interventions generally led to diminished performance in the bonafide class, highlighting the necessity for more advanced methods to accurately model genuine speech. Our findings diverge from those in [30], with EER deteriorating under all asymmetric intervention conditions except for one. This discrepancy is likely due to our one-side distinct strategy of applying interventions exclusively during either the training or testing phase, creating a mismatch. Nevertheless, similar to findings in [30], we also observed that interventions involving loudness were generally less impactful compared to other types of interventions across all evaluations.

### v2

@@ 1 Introduction
Recent progress in voice conversion (VC) and text-to-speech (TTS) technologies have intensified concerns regarding their potential for malicious use, emphasizing the critical role of audio anti-spoofing systems. Audio spoofing detection systems, as binary classifiers, distinguish genuine human speech (bonafide) from artificially generated (spoofed) speech. The main direction in this field is toward advancing the systems’ ability to generalize across unseen spoofing attacks by focusing on learning diverse spoofing artifacts. To this end, a pivotal shift from traditional hand-crafted features [1, 2] to data-driven approaches [3, 4, 5] and data augmentation [6, 7] have been pursued. Research on diversifying training data [8, 9, 10], incorporating domain adaptation strategies [11, 12], and leveraging large-scale pre-trained models [13, 14, 15] also follow this trend.
@@
While in-depth analyses in audio anti-spoofing have significantly advanced the field and provided invaluable insights, several areas remain for further explorations: (i) the majority of studies have focused on spoof class, specifically per-attack interpretations; (ii) analyses on each class exist, yet the emphasis has primarily been on silence, except for [28] exploring other data bias factors; (iii) existing analyses predominantly rely on evaluating outcomes, such as performance metrics and score distributions, without a thorough examination of the internal workings of models during training. Our research diverges from existing studies by delving into the training process through loss analysis and adopting a novel asymmetric intervention on each class and phase(train and test) to understand the respective effects.
Our analysis starts with examining the loss curves for bonafide and spoof classes separately, both with and without the application of RawBoost [7] data augmentation, as illustrated in Figure 1. Note that illustrated loss values are raw values. For the model update, a loss weight of 0.9 (bonafide) and 0.1 (spoof) was employed, considering the imbalanced number of samples in each class. Interestingly, the results indicate that the loss associated with the bonafide class is significantly lower than that for the spoof class. Even if one considers the loss weights in the training phase, a substantial gap still remains between the two classes. These results could signify that the bonafide class is inherently easier to train than the spoof class. Alternatively, this might suggest that modeling the bona fide class is not trivial; however, there exists a shortcut that significantly reduces the training loss. We deploy various loss functions to reveal the meaning behind the low magnitude of the bonafide loss, between the two scenarios mentioned. In particular, the objective functions that can consider the difficulty of samples enable us to understand how the model deals with each class based on our findings.
@@ 3.1 Datasets
ASVspoof2019 [25] is a dataset for logical access (LA) scenario of audio anti-spoofing and it includes 6 and 13 types of spoofing attacks in train/dev and test. The number of utterances for train/dev and test is $50,224$ and $71,237$, respectively.
@@
ASVspoof2021 [37] includes the latest spoofing attacks, which consists of different test sets, each for LA and deepfake (DF) scenarios. The LA and DF subsets include diverse synthetic techniques and audio compressions, respectively.
@@ 3.2 Implementation details
Table 3: Performance comparison in EER when the model is tested on asymmetric ways. All models are trained on the original training dataset but evaluated in different configurations. Here, the ratio indicates a relative change of Te_B compared to that of Te_S.
@@
Table 5: Performance comparison in EER when the model trained on asymmetric ways. All evaluations are conducted using original test set without interventions.
@@ 4.2 Asymmetric test results
Table 3 presents the results of applying asymmetric interventions to a specific class during the test phase. These results are analyzed from two angles: (i) evaluating the intervention’s impact by comparing the original results (O) with those after interventions on the spoof (Te_S) or bonafide (Te_B) classes (columns 2 and 3), and (ii) examining how interventions differently affect the bonafide class compared to the spoof class, as indicated in column 4. From the first angle, enhancements in performance post-intervention hint at possible model biases due to shortcuts. Conversely, declines indicate the intervention’s irrelevance to the target class or might reflect domain differences or unknown factors. The ratio in column 4, calculated as the relative performance impact $(|\textbf{O}-\textbf{Te}\_\textbf{B}|/\textbf{O})/(|\textbf{O}-\textbf{Te}\_\textbf{S}|/\textbf{O})$, offers deeper insight into which class the intervention impacts more. Here, $\text{ratio}>1$ refers to a greater influence on the bonafide class, while $\text{ratio}<1$ implies spoof class is more affected.
Our findings highlight two main observations that align with our analysis of loss functions. First, interventions are prone to improve Te_S (e.g., AASIST with MP3 or AASIST-L with Noise on Te_S), while consistently lowering Te_B performance. This implies that factors leading to quality degradation in utterances tend to be linked with the spoof class, helping the model to classify such inputs as spoofed. The drop in Te_B performance may result from interventions causing bonafide utterances to resemble spoofed ones more closely. Second, the ratio always surpasses 1, as a universal trend regardless of which intervention applied. This emphasizes the bonafide class’s modeling fragility, suggesting that future research should more focus on the robust modeling of the bonafide class.
@@ 4.3 Further test interventions with silence trimming
As introduced in Section 1, the silence could significantly influence the class-wise imbalance results. This prompts an important question: “Do our observations remain valid when silence is excluded from the training or testing datasets or both?” To address this, we conducted experiments with silence removed – referred to as silence trimming – to assess its effect on our findings and show the results in  Table 4. The results show a remarkable consistency across most of our results. When silence was eliminated from both phases, effectively eliminating its influence, the ratio consistently exceeded 1, reaching a peak of 11.23. In additional experiments where silence was removed from only one phase, there were a few instances where the ratio fell below 1; however, the majority still exhibited ratios significantly greater than 1. These consistent outcomes robustly support the conclusion that the presence of silence does not skew our findings.
@@ 4.4 Asymmetric training results
Lastly, we focus on asymmetric interventions during the training phase to explore two key questions: (i) “What happens when interventions are applied solely during training?” and (ii) “Does the class-specific effect of interventions remain consistent if the test set remains untouched?” [^2] The second question is related to the analysis in Section 4.2, where we want to clarify whether performance degradation stems from suddenly appearing unseen domains in the test phase. According to the findings presented in Table 5, an improvement was observed in only one out of nine instances for the Tr_S condition. In contrast, interventions generally led to diminished performance in the bonafide class, highlighting the necessity for more advanced methods to accurately model genuine speech. Please note that we aim to compare the value between Tr_B and Tr_S for each condition to understand how each class is affected by each intervention, not the improvement from the baselines in the first two rows. Our findings diverge from those in [28], with EER deteriorating under all asymmetric intervention conditions except for one. This discrepancy is likely due to our one-side distinct strategy of applying interventions exclusively during either the training or testing phase, creating a mismatch. Nevertheless, similar to findings in [28], we also observed that interventions involving loudness were generally less impactful compared to other types of interventions across all evaluations.
@@ 5 Conclusion
6 Acknowledgment
The work has been partially supported by the Academy of Finland (Decision No. 349605, project “SPEECHFAKES”). Experiments of this work used the Bridges2 system at PSC and Delta system at NCSA through allocations CIS210014 and IRI120008P from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, supported by National Science Foundation grants #2138259,#2138286, #2138307, #2137603, #2138296.